OpenStudy (anonymous):

solve: cos2x+cosx =0 for 0

OpenStudy (anonymous):

Is that a cos squared of x or a cos of twox?

OpenStudy (anonymous):

\[\cos2x=2(cosx)^{2}-1\] use it to make a quadratic in cosx and then solve

OpenStudy (mathmale):

abhi: going to try rishabhjaiswal's suggestion? It's a very good one.

OpenStudy (anonymous):

and how to make a quadratic in cosx

OpenStudy (mathmale):

You're starting with cos2x+cosx =0, or (cos2x)+cosx =0. I put the cos 2x term inside parentheses because we're going to make a substitution, the one suggested by rishabhjaiswal. (See above)

OpenStudy (mathmale):

We get, by substitution, \[(2(\cos x)^{2}-1)+ \cos x=0.\]

OpenStudy (mathmale):

Removing the outer parentheses:\[2(\cos x)^{2}-1+(\cos x)=0.\]

OpenStudy (mathmale):

Rearranging, \[2(\cos x)^{2}+(\cos x) - 1 = 0.\]

OpenStudy (mathmale):

Note that this looks very much like the quadratic equation 2a^2 + a - 1 = 0. Please solve this quadratic equation in a. (We'll return to the previous equation, in (cos x), later.)

OpenStudy (anonymous):

this help me a lot thank you